OpenStudy (anonymous):

is 5(x+9) the same as 5 x+9 ?

hartnn (hartnn):

no

hartnn (hartnn):

a(b+c) = ab+ac

hartnn (hartnn):

thats distributive property

hartnn (hartnn):

5(x+9) = 5*x+5*9 = ... ?

OpenStudy (anonymous):

no dear. the number outside bracket will be multiplied to all numbers inside.

OpenStudy (anonymous):

no is like if its solved would it equal the same?

hartnn (hartnn):

no...

OpenStudy (hba):

Not at all @luxdarlin As @hartnn Showed you.

hartnn (hartnn):

5(x+9) = 5*x+5*9 = 5x+45.

OpenStudy (hba):